OpenStudy (anonymous):

Josh marked 12% of the e-mails in his Inbox as Important. If he has 125 e-mails in his Inbox, how many did he mark as Important?

OpenStudy (1645323):

15

OpenStudy (ddcamp):

12% means that for every 100, there are 12 marked. This also means that for every 1, there are 12/100 marked. There are 125 total, so we have: 125 * (12 / 100) = 125 * 0.12 = 15

OpenStudy (anonymous):

thanks!